Rainbow Pipeline Incident

Clean-up UpdatesSubscribe to Plains Midstream Clean Up Updates

Updates (in reverse chronological order)

  • Fri, 1 February 2013 13:30:00 PM MST

    Site update: Northwest aerial view of ponds and right of way on the remediated footprint.

    Reclamation progress continues at clean-up site

    Ongoing inspections show that reclamation continues to progress well. Aerial patrols and on-the-ground monitoring and testing of the remediated release footprint are ongoing and will continue throughout the winter months.

    Plains will continue to oversee and monitor reclamation activities on site as per its reclamation plan. Plains remains committed to completing a comprehensive site clean up, and will ensure the land is restored to meet applicable environmental standards.
     

    Updated photos and videos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 11 January 2013 13:30:00 PM MST

    Site update: Aerial view of ponds and right of way on the remediated footprint.

    Reclamation progress continues at the site
    The remediated footprint continues to return to conditions similar to the surrounding environment.

    Ongoing inspections show that reclamation is progressing well. Work was completed in December for final touch-up contouring and seeding along the pipeline right-of-way. Aerial and on-the-ground monitoring and testing of the remediated release footprint continues.

    Plains will continue to oversee reclamation activities on site as per its reclamation plan. Plains remains committed to completing a comprehensive site clean up, and will ensure the land is restored to meet applicable environmental standards.

    The next update will be issued Friday, February 1, 2013. Updated photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 7 December 2012 13:30:00 PM MST

    Site update: Aerial view of ponds and right of way on the remediated footprint.

    Reclamation progress continues at clean-up site

    In the nineteen months since the release, Plains has completed remediation activities and initiated reclamation of the release footprint. Ongoing inspections of the footprint show that reclamation is progressing well. 

    Updated photos and videos of the site from October 15 show clear and clean surface water in the ponds and the state of re-growth of the vegetation throughout the remediated release footprint. The 10,000 white spruce seedlings planted within the release footprint and the diversity of plant species on site, shows that vegetation appears to have taken hold. Areas of natural encroachment of native grasses, forbs, shrubs and trees were observed within the footprint as the area returns to conditions similar to the surrounding environment.

    Aerial and on-the-ground monitoring and testing of the remediated release footprint is ongoing. Plains will continue to oversee reclamation activities on site as per its reclamation plan. In the coming years Plains will be providing additional tree plantings as we continue to monitor and re-establish vegetation across the footprint. Plains remains committed to completing a comprehensive site clean up, and will ensure the land is restored to meet applicable environmental standards.

    The next update will be issued Friday, January 4, 2013. Updated photos and videos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 2 November 2012 13:30:00 PM MST

    Site update: The site of the former Command Centre into the first year of reclamation

    Reclamation progress continues at clean-up site

    In the eighteen months since the release, Plains has completed remediation activities and initiated reclamation of the release footprint. Recent inspections of the site have confirmed that the pond water remains clear, and vegetation is re-establishing well. The 10,000 white spruce seedlings planted within the release footprint appear have taken hold and areas of natural encroachment of native grasses, forbes, shrubs and trees were observed. The area continues to return to conditions similar to the surrounding environment.

    Plains will continue to oversee reclamation activities on site as per its reclamation plan. Aerial and on-the-ground monitoring and testing of the remediated release footprint is ongoing. Updated photos and videos of the site from October 15 show the condition of the ponds and the re-growth of the vegetation throughout the remediated release footprint.


    Plains remains committed to completing a comprehensive site clean up, and will ensure the land is restored to meet applicable environmental standards.

    The next update will be issued Friday, December 7, 2012. Updated photos and videos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage

    Communications update:

    • The Rainbow Pipeline incident webpage includes updated photos and videos showing the remediated footprint and current conditions at the site.
    • Plains has initiated phone calls and continues to remain in contact with local First Nations leaders, staff and legal representatives and provided several offers of a MP188 follow up site visit. Where aligned with Plains’ community donations policies, community investment support was provided in October.
    • The Clean-up Updates page of the Rainbow pipeline incident webpage includes an RSS feed, and we encourage interested stakeholders to subscribe for regular updates.

    If you have further questions, please call: 1-866-754-7422 or email: rainbow.incident@plainsmidstream.com.

  • Fri, 5 October 2012 13:30:00 PM MST

    Site update: 

    Reclamation progress continues at clean-up site

    Nearly one year after remediation activities at the clean-up site were completed, restoration continues to progress well. The area continues to return to conditions similar to the surrounding environment. The site continues to re-vegetate to native species of grasses, forbs, shrubs, and trees. Favourable weather and growing conditions have permitted the re-growth of vegetation cover, restoring habitat for area wildlife.

    Ongoing site inspections and assessments confirm that oil contamination has been removed from the site. Remediation of the site is complete, and Plains continues to oversee reclamation activities on site as per its reclamation plan. Aerial and on-the-ground monitoring and testing of the remediated release footprint is ongoing. Updated photos of the site from September 24 show the condition of the ponds and the regrowth throughout the remediated release footprint.

    Plains remains committed to completing a comprehensive site cleanup, and will ensure the land is properly restored to meet applicable environmental standards.

    The next update will be issued Friday, November 2, 2012 Updated photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 7 September 2012 13:30:00 PM MST

    Site update: 

    Reclamation progresses at clean-up site

    As shown in updated photos of the remediated footprint, the area continues to return to conditions similar to the surrounding environment. The site continues to re-vegetate to native species of grasses, forbs, shrubs, and trees. Favourable summer weather and growing conditions have permitted the re-growth of vegetation cover, restoring habitat for area wildlife.

    Ongoing site inspections and assessments confirm that oil contamination has been removed from the site. Remediation of the site is complete, and Plains continues to oversee reclamation activities on site as per its reclamation plan. Aerial and on-the-ground monitoring and testing of the remediated release footprint is ongoing. Updated photos of the site from August 31 show the condition of the ponds and the regrowth throughout the remediated release footprint.

    Plains remains committed to completing a comprehensive site cleanup, and will ensure the land is properly restored to meet applicable environmental standards.

    The next update will be issued Friday, October 5, 2012. Updated photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 3 August 2012 13:30:00 PM MST

    Remediated pond area (July 22)

    Site update:

    Clean-up site progress: Vegetation re-growing, wildlife returning

    The remediated footprint continues to return to conditions similar to the surrounding environment, as the site has started to re-vegetate to native species of grasses, forbs, shrubs, and trees. Local weather conditions have permitted the re-growth of vegetation cover, restoring habitat for area wildlife.

    Recent site inspections and assessments by operations staff and third-party remediation and reclamation specialists have confirmed that oil contamination has been removed from the site. Remediation of the site is complete, and Plains continues to oversee reclamation activities on site as per its reclamation plan. Aerial and on-the-ground monitoring and testing of the remediated release footprint is ongoing. Updated photos of the site from July 20 show the condition of the ponds and the regrowth throughout the remediated release footprint.

    Plains remains committed to completing a comprehensive site cleanup, and will ensure the land is properly restored to meet applicable environmental standards.

    The next update will be issued Friday, September 7, 2012. Updated photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Thurs, 5 July 2012 13:30:00 PM MST

    Site update:

    Reclamation continues at clean-up site

    Plains completed all of the remediation work and much of the reclamation work before winter, and time is now needed for the re-growth of vegetation cover. To date, reclamation work has focused on creating the conditions to allow the remediated footprint to return to conditions similar to the surrounding environment.  As shown in the photos on the Photo Updates page on the Rainbow Pipeline Incident webpage, diverse wildlife has returned to access the ponds within the remediated release footprint.

    Plains continues to oversee on-site reclamation activities. Those activities occur intermittently, and include:

    • inspections of the release footprint;
    • surface and groundwater sampling and analysis;
    • final grading and contouring of the site, if necessary;
    • touch-up seeding and seedling planting to re-establish vegetative cover; and
    • promoting good growing conditions so that the vegetation can re-establish within the cleaned-up areas.

    Plains remains committed to completing the comprehensive site cleanup, and will ensure the land is properly restored to meet all applicable environmental standards.

    The next update will be issued Friday, August 3, 2012. Updated photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 1 June 2012 13:30:00 MST

    Site Update:

    Plains nears completion of a comprehensive site clean up

    Plains completed all of the remediation work and much of the reclamation work before winter, and time is now needed for the natural re-growth of vegetation cover. To date, reclamation work has focused on creating the conditions to allow the remediated footprint to return to conditions similar to the surrounding environment.

    As the ground conditions continue to improve, on-site reclamation activities will resume and focus on:

    • inspections of the release footprint;
    • surface and groundwater sampling and analysis;
    • final grading and contouring of the site, if necessary;
    • touch-up seeding and seedling planting to re-establish vegetative cover; and
    • promoting good growing conditions so that the vegetation can re-establish within the cleaned-up areas. 

    Plains will continue to oversee reclamation activities on site until vegetation cover has been adequately re-established.

    Plains remains committed to completing the comprehensive site cleanup, and will ensure the land is properly restored to meet all applicable environmental standards.

    The next update will be issued Thursday, July 5, 2012. Updated photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 4 May 2012 13:30:00 MST

    Aerial photo of remediated footprint

    Site update:
    Plains remains committed to completing the comprehensive site cleanup, and will ensure the land is properly restored to meet all applicable environmental standards.

     

    As the weather improves and ground conditions allow, on-site activities will resume and focus on:

    • inspections of the release footprint;
    • surface and groundwater sampling and analysis;
    • final grading and contouring of the site, if necessary; and
    • touch-up seeding and seedling planting to re-establish vegetative cover.

    While all of the remediation work and much of the reclamation work was completed before winter, time is now needed for the re-growth of vegetation cover. To date, reclamation work has focused on creating the conditions to allow the remediated footprint to return to conditions similar to the surrounding environment.  The remaining site work will focus on promoting good growing conditions so that the vegetation can re-establish within the clean up areas.
     
    Environmental specialists will monitor the site during the spring transition, and continue with reclamation activities as outlined in the reclamation plan. This monitoring will include collection of surface water samples, collection of groundwater samples, and inspections of the remediated footprint. Inspections will identify any settlement, vegetation or weed issues so they can be addressed. Plains will continue to oversee reclamation activities on site until vegetation cover has been adequately re-established.

    The next update will be issued Friday, June 1, 2012, given the limited activity during spring thaw. Updates on our clean-up operations and photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.      

  • Thu, 5 Apr 2012 13:30:00 MST

    Aerial photo of remediated footprint

    On April 29, 2011, a release of crude oil on Plains’ Rainbow Pipeline was reported about 100 km east of Peace River, Alberta. The released crude oil affected about 8.3 ha (20.5 acres) on the pipeline’s right of way and adjacent treed areas and beaver ponds. More than 49% of available surface crude oil was recovered from the release site within three months of the release. The remaining oil volumes in the soil have been removed from the site and disposed of in approved waste disposal facilities.

    Since completing the remediation of the site in December, the site has remained undisturbed as we wait for spring conditions that will bring regrowth of the reclaimed areas. 

    As warmer weather sets in and ground conditions allow, on-site activities will resume and focus on:

    • inspections of the release footprint;
    • surface and groundwater sampling and analysis;
    • final grading and contouring of the site, if necessary; and
    • touch-up seeding and seedling planting to re-establish vegetative cover.

    Environmental specialists will monitor the site during the spring transition, and continue with reclamation activities as outlined in the reclamation plan. This monitoring will include collection of surface water samples, collection of groundwater samples, and inspections of the remediated footprint. Inspections will identify any settlement, vegetation or weed issues so they can be addressed. Plains will continue to oversee reclamation activities on site until vegetation cover has been adequately re-established.

    The above photo shows the current snow-covered condition of the remediated footprint.

    The next update will be issued Friday, May 4, 2012, given the limited activity during spring thaw. Updates on our clean-up operations and photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.     

  • Fri, 2 Mar 2012 13:30:00 MST

    Site update:

    • Monitoring: Aerial patrols and monitoring of the site are ongoing and will continue throughout the winter.  
    • Access: Rig mat access roads were removed to allow further re-growth to remediated and reclaimed areas in spring; site access is limited to foot travel.
    • Removal of oil-contaminated soil was completed in December. Any final reclamation work can begin once conditions allow in late spring.
    • Reclamation: Significant progress on the reclamation activities have allowed the majority of the reclamation work to be completed before the site was demobilized for the winter. In remediated areas where oil contamination has been removed, clean organic materials previously collected from areas have been placed back to encourage wildlife habitat and vegetation re-growth.
    • In late spring when ground conditions allow, site activities will resume and focus on:
      • inspections of the release footprint;
      • surface and groundwater sampling and analysis;
      • final grading and contouring, if necessary; and
      • touch-up seeding and seedling planting to re-establish vegetative cover.
    • Plains is responsible for the clean-up and will continue to work with government regulatory agencies to ensure that the site is fully remediated and meets all applicable environmental standards.

    The next update will be issued Thursday, April 5, 2012. Updates on our clean-up operations and photos of the site can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.       

  • Fri, 3 Feb 2012 13:30:00 MST

    Site update:

    • We’ve completed work on site for the winter and have demobilized all equipment and personnel from site.
    • Access: Rig mat access roads were removed for the winter, preventing disturbance to the remediated and reclaimed areas and allowing further re-growth in spring.
    • Remediation: Soil remediation was completed in December, and oil-contaminated soil was removed to approved waste disposal facilities.
    • Reclamation: Significant progress on the reclamation activities have allowed the majority of the reclamation work to be completed before the site was demobilized for the winter. In remediated areas where oil contamination has been removed, clean organic materials previously collected from areas have been placed back to encourage wildlife habitat and vegetation re-growth.
    • Monitoring: Aerial patrols and monitoring of the site are ongoing and will continue throughout the winter.  
    • Communications: Additional photos showing the progress of clean-up and remediation activities at the site have been added to the photos webpage.

     

    Safety update:

    • Throughout clean-up, remediation and reclamation activities, we’ve maintained a sharp focus on the Rainbow pipeline clean-up site, and have successfully maintained a strong safety record. We aim to continue this safety record when reclamation operations resume.

     The next update will be issued Friday, March 2, 2012. Updates on our clean-up operations can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 6 Jan 2012 13:30:00 MST

    Site update:

    • Work on site has been completed for the winter and all equipment and personnel have been demobilized from site.
    • Access: Workers have deconstructed all rig mat access roads, and removed all rig mats from the site.  
    • Remediation: Soil remediation has been completed and oil-contaminated soil has been removed to approved waste disposal facilities.
    • Reclamation: Significant progress on the reclamation activities have allowed the majority of the reclamation work to be completed. In remediated areas where oil contamination has been removed, clean organic materials previously collected from areas has been placed back to encourage wildlife habitat and vegetation re-growth. Infrastructure and equipment have been demobilized from the site over winter, preventing disturbance to the remediated and reclaimed areas.
    • Monitoring: Aerial patrols and monitoring of the site are ongoing and will continue throughout the winter.

    Safety update:

    • After thirty-six weeks of mobilization, construction, containment and clean-up operations, the Rainbow pipeline clean-up site has had a sharp focus on safety and has successfully maintained a strong safety record. We aim to continue this safety record when reclamation operations resume.

    The next update will be issued Friday, February 3, 2012. Updates on our clean-up operations can be found on Plains Midstream Canada’s Rainbow Pipeline Incident webpage.

  • Fri, 2 Dec 2011 13:30:00 MST

     

    Site update:

     

    • Work on site has been completed for the winter and all equipment and personnel have been demobilized from site.
    • Access: Workers have deconstructed all rig mat access roads, and have stacked the remaining mats for removal from the site.
    • Remediation: Soil remediation has been completed and the final loads of oil-contaminated soil have been removed to approved waste disposal facilities.
    • Reclamation: Significant progress on the reclamation activities have allowed the majority of the reclamation work to be completed. In remediated areas where oil contamination has been removed, clean organic materials previously collected from areas has been placed back to encourage wildlife habitat and vegetation re-growth. Infrastructure and equipment have been demobilized from the site over winter, preventing disturbance to the remediated and reclaimed areas.
    • Monitoring: Aerial patrols and monitoring of the site are ongoing and will continue throughout the winter.

     

    Safety update:

    • After thirty-one weeks of mobilization, construction, containment and clean-up operations, the Rainbow pipeline clean-up site has had a sharp focus on safety and has successfully maintained a strong safety record.

    The next updates will be issued monthly commencing Friday, January 6, 2012, given the limited activity during the winter months.

  • Fri, 25 Nov 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 38 workers throughout the week. Over the next week, further reductions of personnel on site are anticipated, as the site moves towards demobilization over the winter.
    • Access: Workers continue to deconstruct and remove all rig mat roads that are no longer required in remediated areas. Removal of the access roads is 90% completed.
    • Remediation: Soil remediation is nearing completion and the final loads of oil-contaminated soil will be removed by the end of November.
    • Reclamation: Significant progress on the reclamation activities have allowed the majority of the reclamation work to be completed. In remediated areas where oil contamination has been removed, clean organic materials previously collected from areas are being placed back to encourage wildlife habitat and vegetation re-growth. As part of the reclamation phase, infrastructure and equipment are being demobilized from the site, preventing disturbance to the remediated areas.
    • Monitoring: Aerial patrols and monitoring of the site are ongoing and will continue throughout the winter.

    Safety update:

    • After thirty weeks of mobilization, construction, containment and clean-up operations and over 245,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.

    The next update will be issued on Friday, December 2, 2011. Updates on our clean-up operations can be found on Plains Midstream Canada's Rainbow Pipeline Incident webpage.

     

  • Fri, 18 Nov 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 49 workers throughout the week. Over the next couple of weeks, further reductions of personnel on site are anticipated, as the site moves towards demobilization over winter.
    • Access: Workers continue to deconstruct and remove all rig mat roads that are no longer required in remediated areas.
    • Remediation: Soil remediation is nearing final completion with few remaining stockpiles of oil-contaminated soil requiring removal to approved waste disposal facilities. As the remediation is nearly complete, equipment and personnel are being demobilized from the site.
    • Monitoring: Environmental consultants continue to regularly sample soil stockpiles and monitor water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: Significant progress on the reclamation activities have allowed the majority of the reclamation work to be completed. In remediated areas where oil contamination has been removed along the right-of-way, clean organic materials previously collected from areas are being placed back to encourage wildlife habitat and vegetation re-growth. The pond areas are fully remediated and reclaimed.

    Safety update:

    • After twenty-nine weeks of mobilization, construction, containment and clean-up operations and over 225,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thur, 10 Nov 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 61 workers throughout the week.
    • Access: In an effort to reduce our footprint, workers have deconstructed and removed all rig mat roads that are no longer required in remediated areas.
    • Remediation: Oil-contaminated soil under the rig mat roads has been fully reclaimed. Soil remediation is ongoing, as the remaining stockpiles of oil-contaminated soil continue to be removed to approved waste disposal facilities. The treatment of soil onsite is completed and the Thermal Desorption Unit for treating soil has been demobilized and removed off-site. Over the next few weeks, as the remediation phase is completed, volumes of soil transported from the site will further decrease. Equipment and personnel are gradually demobilizing for the winter.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed along the right-of-way, clean organic materials previously collected from areas are being placed back to encourage wildlife habitat and vegetation re-growth. The pond areas are fully remediated and reclaimed.

    Safety update:

    • After twenty-eight weeks of mobilization, construction, containment and clean-up operations and over 225,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 04 Nov 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 102 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities. In an effort to reduce our footprint, workers are preparing to remove the rig mats that are no longer required, in a staged approach to allow for reclamation in remediated areas.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite has reduced the number of truck trips required to haul soil. Over the next few weeks, as the remediation phase is completed, volumes of soil transported from the site will further decrease. Equipment and personnel will gradually begin to demobilize for the winter.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed, including the pond shorelines and along beaver runs, clean organic materials previously collected from the pond areas are being placed back to encourage wildlife habitat and vegetation re-growth.

    Safety update:

    • After twenty-seven weeks of mobilization, construction, containment and clean-up operations and over 225,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 28 Oct 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 103 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities. In an effort to reduce our footprint, workers are preparing to remove the rig mats that are no longer required, in a staged approach to allow for reclamation in remediated areas.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite has reduced the number of truck trips required to haul soil. Over the next few weeks, as the remediation phase is completed, volumes of soil transported from the site will decrease. Equipment and personnel will gradually begin to demobilize for the winter.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed, including the pond shorelines and along beaver runs, clean organic materials previously collected from the pond areas are being placed back to encourage wildlife habitat and vegetation re-growth.

    Safety update:

    • After twenty-six weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 21 Oct 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 108 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities. In an effort to reduce our footprint, workers are removing access roads and rig mats to allow reclamation in remediated areas.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite has reduced the number of truck trips required to haul soil to approved waste disposal facilities. As the remediation phase is completed, volumes of soil transported from the site will decrease and equipment and personnel will gradually be demobilized for the winter.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed, including the pond shorelines and along beaver runs, clean organic materials previously collected from the pond areas are being placed back to encourage wildlife habitat and vegetation re-growth.

    Safety update:

    • After twenty-five weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 14 Oct 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 113 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities. In an effort to reduce our footprint, workers have relocated the command centre to the north end of the site. The reduction of the rig mat roads on the south access areas allows for the soil to be reclaimed.
    • Remediation: Tarps are being used to cover soil stockpiles at the site to keep soil dry. Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, clean organic materials previously collected from the pond areas are being placed back to encourage wildlife habitat and vegetation re-growth.

    Safety update:

    • After twenty-four weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 07 Oct 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 102 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities. In an effort to reduce our footprint, workers are relocating the command centre to the north end of the site. The reduction of the rig mat roads on the south access areas allows for the soil to be reclaimed.
    • Remediation: Tarps are being used to cover soil stockpiles at the site to keep soil dry. Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation, grasses and small shrubs have begun to return. Clean organic materials previously collected from the pond areas are being placed back to encourage wildlife habitat and vegetation re-growth.

    Safety update:

    • After twenty-three weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 30 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 82 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities. In an effort to reduce our footprint, workers are relocating the command centre to the north end of the site. The reduction of the rig mat roads on the south access areas, allows for the soil to be reclaimed.
    • Remediation: Tarps are being used to cover soil stockpiles at the site to keep soil dry. Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation, grasses and small shrubs have begun to return. Clean organic materials previously collected from the pond areas are being placed back to encourage wildlife habitat and vegetation re-growth.

    Safety update:

    • After twenty-two weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 23 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 78 workers throughout the week.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Water management: Rain earlier in the week required workers to manage rainwater on site, removing released fluids from around the soil stockpiles using vacuum trucks.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities. Tarps are being used to cover soil stockpiles at the site to keep soil dry.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to the pond areas and where rig mats have been removed.

    Safety update:

    • After twenty-one weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Fri, 16 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 92 workers yesterday.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities. Tarps are being used to cover soil stockpiles at the site to keep soil dry.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to the pond areas and where rig mats have been removed.

    Safety update:

    • After twenty weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.

    Communication update:

    • As the majority of the remaining clean up operations continues to focus on soil remediation, Plains will begin issuing weekly updates to highlight the progress. The final daily update will be issued today, and weekly updates will commence each Friday thereafter. The website will also be updated with the weekly clean up report.
  • Thu, 15 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 92 workers yesterday.
    • Access: Workers continue to repair and maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil continue to be dried and removed to approved waste disposal facilities or treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities. Tarps are being used to cover soil stockpiles at the site to keep soil dry.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to the pond areas and where rig mats have been removed.

    Communication update:

    • As the majority of the remaining clean up operations continues to focus on soil remediation, Plains will begin issuing weekly updates to highlight the progress. The final daily update will be issued on Friday, September 16th and weekly updates will commence each Friday thereafter. The website will also be updated with the weekly clean up report.
  • Wed, 14 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 89 workers yesterday.
    • Access: Workers continue to repair maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities or being treated on-site. Treating soil onsite will reduce the number of truck trips required to haul soil to approved waste disposal facilities. Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed and in the pond areas.

    Communication update:

    • As the majority of the remaining clean up operations continues to focus on soil remediation, Plains will begin issuing weekly updates to highlight the progress. The final daily update will be issued on Friday, September 16th and weekly updates will commence each Friday thereafter. The website will also be updated with the weekly clean up report.
  • Mon, 12 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 80 workers over the weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities or being treated on-site.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.
    • Decontamination: We are working to clean equipment and machinery used in the site cleanup that are no longer needed in the current phase of the operation. All equipment and machinery used in the response will be cleaned before it leaves the site.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Fri, 9 Sep 2011 13:30:00 MST

    Site update:

    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities or being treated on-site.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.
    • Decontamination: We are working to clean equipment and machinery used in the site cleanup that are no longer needed in the current phase of the operation. All equipment and machinery used in the response will be cleaned before it leaves the site.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Thu, 8 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 99 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling and clean-up activities.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities or being treated on-site. Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Small backhoes are being used to remove remaining pockets of oil from the pond shorelines and islands.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.
    • Decontamination: We are working to clean equipment and machinery used in the site cleanup that are no longer needed in the current phase of the operation. All equipment and machinery used in the response will be cleaned before it leaves the site.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Wed, 7 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 90 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities or being treated on-site. Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Small backhoes are being used to remove remaining pockets of oil from the pond shorelines and islands.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.
  • Tue, 6 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 77 workers over the long weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities or being treated on-site. Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Small backhoes are being used to remove remaining pockets of oil from the pond shorelines and islands.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Fri, 2 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 93 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Soil remediation is ongoing, as stockpiles of oil-contaminated soil are being removed to approved waste disposal facilities and readied for on-site treatment. Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Small backhoes are being used to remove remaining pockets of oil from the pond shorelines and islands. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.

    Safety Update

    • After eighteen weeks of mobilization, construction, containment and clean-up operations and over 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thu, 1 Sep 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 87 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In remediated areas where oil contamination has been removed including the pond shorelines and along beaver runs, native species of vegetation have shown re-growth. Grasses and small shrubs have also begun returning to areas where rig mats have been removed.
  • Wed, 31 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 92 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Water management: Crews have focused on managing rainwater on site, monitoring water levels in the ponds and removing released fluids from around the soil stockpiles using vacuum trucks.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: Although late in the 2011 growing season, native species of vegetation have shown re-growth on the pond shorelines, along beaver runs, and in other remediated areas where oil contamination has been removed. Grasses and small shrubs have begun returning to areas where timber rig mats have been removed.
  • Tue, 30 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 71 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Water management: Crews have focused on managing rainwater on site, monitoring water levels in the ponds and removing released fluids from around the soil stockpiles using vacuum trucks.
    • Monitoring: Environmental consultants regularly sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: Although late in the 2011 growing season, native species of vegetation have shown re-growth on the pond shorelines, along beaver runs, and in other remediated areas where oil contamination has been removed. Grasses and small shrubs have begun returning to areas where timber rig mats have been removed.

    Operation update:

    • Plains commenced the restart of the NPS 20" Rainbow Pipeline this morning. Crude oil flows on the pipeline, which transports crude oil from Zama to Nipisi, AB, will be restored gradually over the next seven days.
  • Mon, 29 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 85 workers over the weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In the remediated areas where oil contamination has been removed, clean woody debris has been spread around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Operation update:

    • On Friday, August 26, 2011, Plains Midstream Canada (Plains) received approval from the ERCB to restart the NPS 20" Rainbow Pipeline. Plains intends to resume service on the northern segment of the pipeline on Tuesday, August 30, 2011. The line has been shut down since a crude oil release was confirmed on Friday, April 29, 2011.
    • Crude oil flows on the pipeline, which transports crude oil from Zama to Nipisi, AB, will be restored gradually over a seven-day period following initial restart.
  • Fri, 26 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 79 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In the remediated areas where oil contamination has been removed, clean woody debris has been spread around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Safety update:

    • After seventeen weeks of mobilization, construction, containment and clean-up operations and approximately 200,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.

    Operation update:

    • Plains Midstream Canada (Plains) has received approval from the ERCB to restart the NPS 20" Rainbow Pipeline. Plains intends to resume service on the northern segment of the pipeline on Tuesday, August 30, 2011. The line has been shut down since a crude oil release was confirmed on Friday, April 29, 2011.
    • Crude oil flows on the pipeline, which transports crude oil from Zama to Nipisi, AB, will be restored gradually over a seven-day period following initial restart.
  • Thu, 25 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 97 workers over the weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated. We're estimating the clean-up and soil remediation phase will take a total of eight months to complete.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.
  • Wed, 24 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 87 workers over the weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Tue, 23 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 86 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches and around soil stockpiles.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.

    Operational update:

    • Plains remains committed to keeping our stakeholders informed of the progress on our clean-up operations. As such, we've attached a three-month summary update on our clean-up operations.
  • Mon, 22 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 70 workers over the weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches and around soil stockpiles.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Fri, 19 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 79 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches and around soil stockpiles.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.

    Safety Update

    • After sixteen weeks of mobilization, construction, containment and clean-up operations, and approximately 195,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.

    The next daily update will be issued Monday afternoon. Regular updates on our clean-up operations can be found on Plains Midstream Canada's Rainbow Pipeline Incident webpage.

  • Thu, 18 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 74 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches and around soil stockpiles.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil-contaminated soil has been removed, the site will be considered 100% remediated.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Environmental Update

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Wed, 17 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 77 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches and around soil stockpiles.
    • Remediation: Workers are removing oil-contaminated soil from the area adjacent to the pipeline right of way. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil contaminated soil has been removed, the site will be considered 100% remediated.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Operations update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Tue, 16 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 80 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns. Once the oil contaminated soil had been removed, the site will be considered 100% remediated.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews have begun spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape. Clean organic material and mulch will be spread over the underlying clay to retain moisture and allow seeds to germinate.

    Operations update:

    • Plains has received conditional approval from the ERCB to resume operations of the northern section of Plains' Rainbow Pipeline, which transports crude oil from Zama to Nipisi, AB. The ERCB has outlined conditions that need to be met before the Rainbow Pipeline can be returned to service. We are currently reviewing the conditions, and will assess when we can resume operations. At this point in time, we do not have a firm timeline for restart of the pipeline.
  • Mon, 15 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 80 workers over the weekend.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews have begun spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Fri, 12 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 67 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling. The location for the on-site soil treatment equipment is being prepared.
    • Remediation: Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews have begun spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation and to allow the site to be re-contoured to match the surrounding landscape.

    Safety update:

    • After fifteen weeks of mobilization, construction, containment and clean-up operations, with approximately 188,000 work hours logged, the Rainbow pipeline clean-up site continues to have a sharp focus on safety and has been successful in maintaining a strong safety record. Daily safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thur, 11 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 71 workers yesterday.
    • Access: Workers continue to maintain the main gravel access road and rig mat roads throughout the clean-up site to allow ongoing soil hauling. The location for the on-site soil treatment equipment is being prepared.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Remediation: Tarps are being used to cover soil stockpiles at the site, to keep soil piles dry and facilitate transport during rainy conditions. Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Reclamation: In the remediated areas where oil contamination has been removed, crews have begun spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation.
  • Wed, 10 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 67 workers yesterday.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at approved waste disposal facilities. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Reclamation: Plains is in the early stages of mobilizing crews and materials. In the remediated areas where oil contamination has been removed, crews are spreading clean woody debris around the ponds to encourage re-growth of native species of vegetation. In addition, coconut fibre "geotextile" stabilization material will be in place to restore habitat. With the growing season coming to a close at MP 188, reclamation crews are readying the site conditions for regrowth in the spring.
  • Tue, 9 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 51 workers yesterday.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
  • Mon, 8 Aug 2011 13:30:00 MST

    Site Update:

    • Daily personnel count on site averaged 50 workers over the weekend.
    • Workers have used vacuum trucks and skimmers to recover nearly all available free oil within the footprint, which equals nearly half of the estimated volume of crude oil released. Oil-contaminated soil has also been scraped from the right-of-way and pond areas, and stockpiled for transport and proposed on-site treatment.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Over the weekend, crews focused their efforts on managing rainwater on site, monitoring water levels in the ponds and removing release fluids from containment trenches and around the soil stockpiles using vacuum trucks.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: With the completion of oil and oil-contaminated soil removal from the ponds, Plains is moving towards commencing reclamation work.
  • Fri, 5 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 61 workers yesterday.
    • Workers have used vacuum trucks and skimmers to recover nearly all available free oil within the footprint, which equals nearly half of the estimated volume of crude oil released. Oil-contaminated soil has also been scraped from the right-of-way and pond areas, and stockpiled for transport and proposed on-site treatment.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Oil recovery: Vacuum trucks continue to remove release fluids from containment trenches along the right of way and around soil stockpiles.
    • Remediation: Stockpiles of oil-impacted soil (scraped from the release footprint) continue to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: With the completion of oil and oil-contaminated soil removal from the ponds, Plains is moving towards commencing reclamation work.

    Safety Update:

    • After fourteen weeks of mobilization, construction, containment and clean-up operations, more than 183,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. Safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thu, 4 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 66 workers yesterday.
    • Workers have used vacuum trucks and skimmers to recover nearly all available free oil within the footprint, which equals nearly half of the estimated volume of crude oil released. Oil-contaminated soil has also been scraped from the right-of-way and pond areas, and stockpiled for transport and proposed on-site treatment. A preliminary balance calculation, which includes the oil entrained in the contaminated soils (from the release footprint), appears to account for all the released oil. Final calculations will only be possible once the line balance has been measured (requiring the pipeline to be reactivated) and once all the oil-contaminated soils have been excavated, homogenized, sampled and analyzed.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Oil recovery: Workers continue to remove oil remnants from sections of the beaver runs. Vacuum trucks continue to remove release fluids from containment trenches along the right of way and around soil stockpiles.
    • Remediation: Stockpiles of oil-impacted soil (scraped from the release footprint) continue to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: With the completion of oil and oil-contaminated soil removal from the ponds, Plains is moving towards commencing reclamation work.
  • Wed, 3 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 61 workers yesterday.
    • As of 7:00 a.m. on August 3rd, 8081 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2221 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Oil removal: Workers continue to remove oil remnants from sections of the beaver runs.
    • Remediation: Stockpiles of oil-impacted soil (scraped from the release footprint) continue to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Monitoring: Environmental consultants continue to sample soil stockpiles and collect water samples from the network of groundwater monitoring wells installed around the release footprint. Test results indicate the release impacts remain well contained.
    • Reclamation: With the completion of oil and oil-contaminated soil removal from the ponds, Plains is moving towards commencing reclamation work.
  • Tue, 2 Aug 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 60 workers yesterday.
    • As of 7:00 a.m. on August 2nd, 8081 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2221 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Remediation: Stockpiles of oil-impacted soil (scraped from the release footprint) continue to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Reclamation: With the completion of oil and oil-contaminated soil removal from the ponds, Plains is moving towards commencing reclamation work.
  • Thu, 28 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 74 workers yesterday.
    • As of 7:00 a.m. on July 28th, 8081 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2221 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Crews continue to focus their efforts on managing rainwater on site, monitoring water levels in the ponds and removing release fluids from containment trenches and around the soil stockpiles using vacuum trucks.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling. Preparations are underway to accommodate transport of on-site soil treatment equipment.
    • Soil remediation: Impacted soil has been removed from all the pond areas. The ponds were dewatered and small backhoes were used to remove remaining pockets of oil from the pond shorelines and islands.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
  • Fri, 29 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 63 workers yesterday.
    • Since the April 29 release from the Rainbow pipeline, Plains has mounted an aggressive clean-up response. Initial efforts focused on access and construction work to gain access to contain and remove oil throughout the site. At the same time as access was being built, large crews of workers were removing bulk surface oil from the right-of-way and pond areas, and working to contain the oil. This intensive, full-scale response helped limit the size of the footprint from the release.
    • In the three months since the release, nearly half (49%) of the original estimated crude oil volume has been recovered. As of July 29th, 2221 m3 of clean oil has been recovered, and a total of 8081 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site.
    • The available free oil within the footprint has almost entirely been recovered, which equals roughly half of the estimated volume of crude oil released. A preliminary balance calculation, which includes the oil entrained in the contaminated soils (from the release footprint), appears to account for all the released oil. A final balance calculation will only be possible once the line balance has been measured (requiring the pipeline to be reactivated) and once all the oil-contaminated soils have been excavated, homogenized, sampled and analyzed.
    • Plains will continue with its efforts to remediate the release footprint. We remain committed to completing a comprehensive clean up and we will continue to work until the land is restored to meet all applicable environmental standards. We will continue to monitor the site for years after.
    • Remediation: Stockpiles of oil-impacted soil (scraped from the release footprint) continue to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.

    Safety Update

    • After thirteen weeks of mobilization, construction, containment and clean-up operations, more than 177,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. Safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.

    Operation Update

    • Following a thorough analysis of the Rainbow Pipeline incident, Plains submitted a request to the ERCB to resume operation of the Rainbow Pipeline.
    • Plains has addressed all requests outlined by the ERCB, and continues to work closely with the ERCB.
    • The pipeline's return to service is contingent on receiving regulatory approval. We do not have a firm timeline for restart of the pipeline.
  • Wed, 27 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 78 workers yesterday.
    • As of 7:00 a.m. on July 26th, 7974 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2216 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Soil remediation: Impacted soil has been removed from all the pond areas. The ponds were dewatered and small backhoes were used to remove remaining pockets of oil from the pond shorelines and islands.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.

    Environmental update:

    • Daily wildlife perimeter fence checks and patrols continue as part of an effort to minimize any additional impacts on wildlife.
  • Tue, 26 Jul 2011 17:16:00 MST

    Site update:

    • Daily personnel count on site totalled 63 workers yesterday.
    • As of 7:00 a.m. on July 26th, 7974 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2216 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Soil remediation: Impacted soil has been removed from all the pond areas. The ponds were dewatered and small backhoes were used to remove remaining pockets of oil from the pond shorelines and islands.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards
  • Mon, 25 Jul 2011 17:16:00 MST

    Site update:

    • Daily personnel count on site totalled 60 workers over the weekend.
    • As of 7:00 a.m. on July 25th, 7974 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2216 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Crews continue to focus their efforts on managing rainwater on site, monitoring water levels in the ponds and removing release fluids from containment trenches and around the soil stockpiles using vacuum trucks.
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site to allow ongoing soil hauling.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
  • Fri, 22 Jul 2011 17:16:00 MST

    Site update:

    • Daily personnel count on site totalled 74 workers yesterday.
    • As of 7:00 a.m. on July 22nd, 7918 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2216 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Remediation: Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • As the final stages of remediation work on the easternmost pond are completed, work will shift from labour-intensive work to the mechanical-intensive work required for final soil remediation and site reclamation, lessening the need for a large on-site workforce. Plains remains committed to local hiring.

    Environmental update:

    • After twelve weeks of mobilization, construction, containment and clean-up operations, more than 170,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. Safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thu, 21 Jul 2011 17:16:00 MST

    Site update:

    • Daily personnel count on site totalled 84 workers yesterday.
    • As of 7:00 a.m. on July 21st, 7918 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2216 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on repairing and maintaining rig mat roads throughout the clean-up site.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Remediation: Soil remediation continues near the easternmost pond, and crews continue flushing and cleaning nearby beaver runs. Samples are taken regularly from the soil stockpiles for analysis to determine the extent of oil contamination. Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility. Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
  • Wed, 20 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totalled 79 workers yesterday.
    • Recovered volumes are unavailable today, but will be included in tomorrow's daily update.
    • Access: Work continues on repairing and maintaining rig mat roads throughout the clean-up site.
    • Oil recovery: Vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Remediation: Soil remediation continues on the right of way and near the easternmost pond area. Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility.
    • Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.

    Environmental update:

    • Daily wildlife perimeter fence checks and patrols continue as part of an effort to minimize any additional impacts on wildlife.
  • Tue, 19 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 77 workers yesterday.
    • As of 7:00 a.m. on July 19th, 7918 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2216 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Workers continue to repair and maintain access roads at the release site.
    • Remediation: In the easternmost pond area, work is 90% compete on removing oil-stained vegetation, focusing on the shoreline and small 'islands' in the pond, cleaning the beaver runs, and removing and replacing oil-impacted sections of the beaver dam. Oil-impacted soil continues to be removed from the site for disposal at an approved waste disposal facility.
    • Additional precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Mon, 18 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 80 workers over the weekend.
    • As of 7:00 a.m. on July 18th, 7910 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2215 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Compacted clay bases, trenching and culverts have been installed under the access road in support of daily operations and to improve water management.
    • Remediation: Work continues on soil remediation in the easternmost pond area, with workers removing oil-stained vegetation, focusing on the shoreline and small 'islands' in the pond. The first loads of oil-impacted soil were removed from the site for disposal at an approved waste disposal facility, and extra precautionary measures have been applied to all loads to prevent odours and mitigate community concerns.
    • Plains has issued a communication bulletin to the local communities and regulatory authorities providing an update on soil handling plans, as previously identified in our waste management plans. Plains remains committed to local hiring and keeping in regular contact with local communities.
  • Fri, 15 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 78 workers yesterday.
    • As of 7:00 a.m. on July 15th, 7910 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2215 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on repairing and maintaining rig mat roads throughout the clean-up site.
    • Oil recovery: In easternmost pond area, workers are removing oily wood and debris.
    • Remediation: Soil remediation continues on the easternmost pond area, and is nearing 85% complete.

    Safety update:

    • After eleven weeks of mobilization, construction, containment and clean-up operations, more than 163,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. Safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thu, 14 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 71 workers yesterday.
    • Site conditions continue to improve, as warmer weather dries the clean-up site.
    • As of 7:00 a.m. on July 14th, 7891 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2214 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on repairing and maintaining rig mat roads throughout the clean-up site.
    • Oil recovery: Crews are removing released fluid from containment trenches, bell holes, and around the soil stockpiles using vacuum trucks. In the easternmost pond area, workers are removing oil-stained wood and vegetation from the pond.
    • Remediation: Soil remediation continues on the easternmost pond area, and is nearing 85% complete. Contaminated soil has been stockpiled in windrows within the release footprint and soil samples have been taken from the soil stockpiles for analysis to determine the extent of oil contamination.

    Environmental update:

    • Daily wildlife perimeter fence checks and patrols are completed to minimize any additional impacts on wildlife.
  • Wed, 13 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 78 workers yesterday.
    • Crews continue to focus their efforts on managing rainwater that has accumulated on site, monitoring water levels in the ponds and removing released fluid from containment trenches and around the soil stockpiles using vacuum trucks.
    • As of 7:00 a.m. on July 12th, 7843 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2211 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on repairing and maintaining rig mat roads throughout the clean-up site.
    • Remediation: Soil remediation continues on the right of way and near the easternmost pond area.
    • Updated site photos have been added to the Photos and Media page on the Rainbow Pipeline Incident webpage.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Tue, 12 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 82 workers yesterday.
    • Crews continue to focus their efforts on managing rainwater that has accumulated on site, monitoring water levels in the ponds and removing released fluid from containment trenches using vacuum trucks. Workers continue to maintain berms around soil piles and ensuring appropriate trenching is in place.
    • As of 7:00 a.m. on July 12th, 7787 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2211 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • As work at the clean-up site shifts from labour-intensive work required for containment, access construction, oil removal, and oil recovery, to the mechanical-intensive work required for soil remediation and site reclamation, the need for a large on-site workforce will gradually lessen. Plains remains committed to local hiring.
  • Mon, 11 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 90 workers over the weekend.
    • Heavy rains over the weekend scaled back operations at the clean-up site. Recent sunny weather has begun to improve conditions on site.
    • As of 7:00 a.m. on July 11th, 7703 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2211 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Crews have focused their efforts on managing rainwater that has accumulated on site, monitoring water levels in the ponds and removing released fluid from containment trenches using vacuum trucks. Workers are maintaining berms around soil piles and ensuring appropriate trenching is in place.
    • Oil recovery: One skimmer was in use over the weekend to recover free product.
    • Since the early days of the response, Plains has worked to ensure that local communities are well represented in the on-site workforce. As the clean up progresses, the large workforce initially required for the access and construction work and manual labour for oil recovery and site cleanup will gradually reduce as the site moves out of the emergency response and oil recovery phase and into the soil remediation and site reclamation phase. Plains remains committed to local hiring.

     

  • Fri, 8 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 105 workers yesterday.
    • As of 7:00 a.m. on July 8th, 7425 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2210 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Oil recovery: Crews continue flushing and cleaning the beaver dam and beaver runs in the central pond area. Processed peat moss, which acts as an oil absorbent, continues to be applied to the banks of the beaver runs to help remove residual oil.
    • Three vacuum trucks and one gator vacuum truck continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Soil remediation continues on the right of way near the central and west pond areas.

    Safety update:

    • After ten weeks of mobilization, construction, containment and clean-up operations, more than 156,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. Safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Thu, 7 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 105 workers yesterday.
    • Personnel: Although staff During the current clean-up site operations, the approximate numbers and different roles on-site typically include:
      • 34 general clean-up and site maintenance workers;
      • 20 contractors responsible for safety, site security, air monitoring, and paramedical support;
      • 7 vacuum truck and gator vacuum truck drivers;
      • 10 heavy equipment operators;
      • 8 environmental and spill-response consultants;
      • Two teams of 2 wildlife patrol team members;
      • Plains employees, and other personnel as needed, including surveyors, welders, pipefitters, and mechanics.
    • As of 7:00 a.m. on July 7th, 7401 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2210 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Oil recovery: Crews continue flushing and cleaning the beaver dam and beaver runs in the central pond area. Processed peat moss, which acts as an oil absorbent, is being applied to the banks of the beaver runs to help remove residual oil.
    • Four vacuum trucks and two gator vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Soil remediation continues on the right of way near the central and west pond areas. Samples have been taken from the soil stockpiles for analysis to determine the extent of oil contamination.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Wed, 6 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 108 workers yesterday.
    • As of 7:00 a.m. on July 6th, 7363 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2210 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Oil recovery: Crews continue flushing and cleaning the beaver dam and beaver runs in the central pond area. Processed peat moss, which acts as an oil absorbent, is being applied to the banks of the beaver runs to help remove residual oil.
    • Four vacuum trucks and two gator vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Soil remediation continues on the right of way near the central pond area; nearly 95% of the contaminated soil has been removed from the right of way. Samples have been taken from the soil stockpiles for analysis to determine the extent of oil contamination.

    Environmental update:

    • Daily wildlife perimeter fence checks and patrols continue as part of an effort to minimize any additional impacts on wildlife. Before dawn and after dusk, air horns are used to keep incoming birds away.
  • Tue, 5 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 97 workers yesterday.
    • Recent warm weather has improved site conditions.
    • As of 7:00 a.m. on July 5th, 7279 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2210 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: New rubber rig mats have been installed on the east road.
    • Oil recovery: Crews continue flushing and cleaning the beaver dam and beaver runs in the central pond area. Processed peat moss, which acts as an oil absorbent, is being applied to the banks of the beaver runs to help remove residual oil. Four vacuum trucks and two gator vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Soil remediation continues on the right of way near the central pond area.
  • Mon, 4 Jul 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 99 workers over the weekend.
    • Site conditions continue to improve, as warm weather dries the clean-up site.
    • As of 7:00 a.m. on July 4th, 7188 m3 of recovered volumes (which includes oil, water and solids) has been collected from the release site, including 2210 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: New rubber rig mats have been installed on site, as well as additional mile markers and signage indicating speed limits.
    • Oil recovery: Crews continue flushing and cleaning the beaver dam and beaver runs in the central pond area. Five vacuum trucks and three gator vacuum trucks continue to remove released fluids from containment trenches along the right of way and around soil stockpiles.
    • Soil remediation continues on the right of way near the central pond area, and is approximately 70% complete.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Thu, 30 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 100 workers yesterday.
    • Site conditions are improving, allowing the site to dry. Full clean-up operations are gradually resuming.
    • As of 7:00 a.m. on June 30th, 6838 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2203 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: New rubber rig mats have been delivered and work continues on maintaining rig mat roads throughout the clean-up site.
    • Oil recovery: Workers are manually removing oil from the central pond using long-handled pan skimmers. Five vacuum trucks and three gator vacuum trucks continue to remove spill fluid from containment trenches along the right of way and around soil stockpiles. Storage cells continued to be constructed for soil stockpiles.

    Safety update:

    • After nine weeks of mobilization, construction, containment and clean-up operations, more than 147,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. Safety briefings are held to identify any potential hazards on site, reiterating known site hazards and to provide general safety awareness.
  • Wed, 29 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 96 workers yesterday.
    • Recent sunny weather has improved conditions on site, allowing the site to dry and full clean-up operations to gradually resume.
    • As of 7:00 a.m. on June 29th, 6706 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2199 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Work continues on maintaining rig mat roads throughout the clean-up site, and completing work on the east right-of-way road and the south access road.
    • Oil recovery: Five vacuum trucks and three gator vacuum trucks continue to remove spill fluid from containment trenches along the right of way and around soil stockpiles.
    • Remediation: The eastern pond has had the majority of surface oil skimmed, and oil-stained vegetation removed from approximately three quarters of its shoreline. The western pond has been completely remediated. Work continues in the central pond area.
  • Tue, 28 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 98 workers yesterday.
    • Recent sunny weather has improved conditions on site, allowing full clean-up operations to gradually resume.
    • As of 7:00 a.m. on June 28th, 6626 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2196 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Crews are maintaining and improving rig mat roads throughout the clean-up site, and will continue to add more wooden and rubber rig mats to improve access.
    • Oil recovery: Seven vacuum trucks and three gator vacuum trucks continue to remove spill fluid from containment trenches along the right of way and around soil stockpiles.
    • Remediation work continues to focus on the central pond area.
  • Mon, 27 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 65 workers over the weekend.
    • While it rained on Saturday, the weather improved on site on Sunday. As conditions dry, the site will fully staff up again.
    • Crews continue to focus their efforts on managing accumulated rainwater on site and monitoring water levels in the ponds.
    • As of 7:00 a.m. on June 27th, 6486 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2196 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Access: Crews are maintaining rig mat roads throughout the clean-up site, and will be adding wooden and rubber rig mats to improve access.
    • Oil recovery: Over the weekend, seven vacuum trucks and three gator vacuum trucks were on site removing spill fluid from containment trenches along the right of way and around soil stockpiles.
    • Remediation work continues on the central pond area.

    Environmental update:

    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Fri, 24 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 85 workers yesterday.
    • Clean-up operations remain scaled back due to wet conditions on the clean-up site. Crews continue to focus their efforts on managing rainwater that has accumulated on site and monitoring water levels in the ponds.
    • As of 7:00 a.m. on June 24th, 6114 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2192 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Oil recovery: Nine vacuum trucks and three gator vacuum trucks are on site removing spill fluid from containment trenches along the right of way and around soil stockpiles.
    • Remediation: The eastern pond has had the majority of surface oil skimmed, and oil-stained vegetation removed from approximately 70% of its shoreline. The western pond has been completely remediated. Work continues on cutting, raking, and flushing the shoreline and beaver runs in the central pond area.

    Safety update:

    • After eight weeks of mobilization, construction, containment and clean-up operations, more than 141,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. We continue to hold daily safety briefings to identify any potential hazards on site, reiterating known site hazards, general safety awareness and recognizing fatigue.
  • Thu, 23 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 89 workers yesterday.
    • Clean-up operations remain scaled back due to wet conditions on the clean-up site. Crews continue to focus their efforts on managing rainwater that has accumulated on site and monitoring water levels in the ponds.
    • As of 7:00 a.m. on June 23rd, 5889 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2191 m3 of clean oil (approximately 49% of the original estimated 4500 m3 released).
    • Oil recovery: Nine vacuum trucks and two gator vacuum trucks are on site removing spill fluid from containment trenches along the right of way and around soil stockpiles.
    • Decontamination: Workers continue to clean equipment and machinery used in the site cleanup that are no longer needed in the current phase of the operation. All equipment and machinery used in the response will be cleaned before it leaves the site.
    • Updated site photos have been added to the Photos and Media page on the Rainbow Pipeline Incident webpage.

     

  • Wed, 22 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 99 workers yesterday.
    • Clean-up operations remain scaled back due to heavy rain and weather-related safety concerns. Once weather improves and conditions allow, plans are to fully staff up again.
    • Due to the scaled-back clean-up operations, the recovered oil volumes were limited. As of 7:00 a.m. on June 22nd, 5475 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2181 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Access: Rig mats near the command centre have been removed to construct storage cells for soil stockpiles.
    • Crews have focused their efforts on managing rainwater that has accumulated on site, monitoring water levels in the ponds and removing spill fluid from containment trenches using gator vacuum trucks.

     

  • Tue, 21 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site totaled 62 workers yesterday.
    • Clean-up operations remained scaled back due to weather-related safety concerns. Once weather improves, plans are to fully staff up again.
    • Access: Timber rig mats on the north end of the clean-up site are being moved to allow the road to be re-built to accommodate heavy equipment. Maintenance continues on segments of road throughout the site, and adjusting the placement of rig mats that have shifted.
    • Due to the scaled-back clean-up operations, the recovered oil volumes were limited. As of 7:00 a.m. on June 21th, 5256 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2178 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Oil recovery: Gator vacuum trucks continue removing collected oil from containment trenches along the right-of-way and around the soil stockpiles.
  • Mon, 20 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site averaged 58 workers over the weekend.
    • Clean-up operations remained scaled back over the weekend due to weather-related safety concerns. Once weather improves, plans are to fully staff up again.
    • Due to the scaled-back clean-up operations, the recovered oil volumes were limited. As of 7:00 a.m. on June 20th, 5194 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2178 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Oil recovery: Gator vacuum trucks were operating to remove collected oil from containment trenches along the right-of-way and around the soil stockpiles. In addition to the mechanical skimmers operating in the pond area, workers are manually removing oil from areas near the booms using long-handled pan skimmers.

    Environmental Update

    • Bears have been sighted near the perimeter of the clean-up area. The Fish and Wildlife division of Alberta Sustainable Resource Development will work to relocate bears away from the site. Garbage bins have been relocated, and additional bear-proof garbage bins have been added.
    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.

     

  • Fri, 17 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 66 workers yesterday.
    • Clean up operations were scaled back due to weather-related safety concerns.
    • Due to the scaled-back clean up operations, the recovered oil volumes were limited. As of 7:00 a.m. on June 17th, 5034 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2172 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Oil recovery: Four vacuum trucks were operating to remove collected oil from containment trenches and bell holes.
    • Plains continues to work closely with government regulatory bodies on clean-up plans for the release site.

    Safety Update

    • After seven weeks of mobilization, construction, containment and clean-up operations, more than 135,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents and continues to have a sharp focus on safety. We continue to hold daily safety briefings to identify any potential hazards on site, reiterating known site hazards and general safety awareness.

     

  • Thu, 16 Jun 2011 12:00:00 MST

    Site update:

    • Daily personnel count on site peaked at 67 workers yesterday.
    • Clean up operations were further scaled back due to the weather conditions and safety concerns related to the continuous rain.
    • Due to the scaled-back clean up operations from the weather conditions, the recovered oil volumes was limited. As of 7:00 a.m. on June 16th, 4925 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2169 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).   
    • Workers, with the use of gator vacuum trucks, continue to recover oil from the collection trenches along the right-of-way and around the soil stockpiles.  
    • Monitoring: Installation of the groundwater monitoring wells has been completed. The long-term groundwater monitoring at the site is a precautionary measure to identify any potential impacts.
    • Environmental Update
    • Release site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.  
  • Wed, 15 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 152 workers yesterday.
    • Clean up operations were scaled back yesterday due to rainy weather conditions and safety concerns related to slipping hazards on the rig mats. The use of heavy and large equipment has also decreased due to access and travel challenges on the rig mats. Oil recovery was suspended as the vacuum trucks were unable to be deployed.
    • Due to the weather conditions on site, the recovered volumes have not changed. As of 7:00 a.m. on June 15th, 4824 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2167 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).  
    • Monitoring: Bore holes have been installed for wells that will be used to monitor any potential oil movement outside of the established perimeter of the release area. Groundwater monitoring wells continue to be installed for long-term groundwater monitoring at the site as a precautionary measure to identify any potential impacts.

    Environmental Update

    • Daily wildlife perimeter fence checks and patrols are completed in order to minimize any additional impacts on wildlife. Air horns are used to keep incoming birds away from before dawn to after dusk.

     

  • Tue, 14 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 158 workers yesterday.
    • As of 7:00 a.m. on June 14th, 4824 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2167 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Oil recovery: : Five skimmers were in operation and eight vacuum trucks were used to remove collected oil from containment trenches and bell holes.
    • In the pond area, workers continue to remove oily, woody debris and to wash the pond’s shoreline and dormant beaver runs with fresh water in order to facilitate oil recovery.
    • Soil remediation: Remediation continues on the top layer of contaminated soil, as it is removed and stockpiled on the north end of the right of way.
    • Monitoring: Bore holes have been installed for wells that will be used to monitor any potential oil movement outside of the established perimeter of the release area. Groundwater monitoring wells have also been installed for long-term groundwater monitoring at the site as a precautionary measure to identify any potential impacts.

     

  • Mon, 13 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 154 workers yesterday.
    • As of 7:00 a.m. on June 13th, 4732 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2156 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Access: Nearly 2 km of timber and composite rig mats were re-positioned over the weekend to allow greater access throughout the clean-up site.
    • Oil recovery: : Over the weekend, an average of 7 skimmers were in daily operation in the pond areas. There's been significant progress in removing surface oil from the ponds, so fewer skimmers are required for the remaining area.
    • In the pond area, workers continue to cut and rake dried seasonal vegetation and remove oily, woody debris. They also continue to wash the pond's shoreline and dormant beaver runs with fresh water in order to facilitate oil recovery.
    • Decontamination: We are working to clean equipment and machinery used in the site cleanup that are no longer needed in the current phase of the operation. All equipment and machinery used in the response will be cleaned before it leaves the site.

    Environmental update:

    • Spill site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Fri, 10 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 138 workers yesterday.
    • As of 7:00 a.m. on June 10th, 4440 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2142 m3 of clean oil (approximately 48% of the original estimated 4500 m3 released).
    • Access: Workers continue to reset the location of timber rigs mats, to accommodate the soil stockpiles and allow the areas previously covered by the mats to undergo soil remediation. Additional mats are being placed to allow greater access throughout the clean-up area.
    • Oil recovery: Workers continue to cut and rake dried seasonal vegetation and wash the pond’s shoreline and dormant beaver runs with fresh water in order to facilitate oil recovery.
    • Workers continue to use vacuum trucks to collect oil from the perimeter trenches, bell holes, and collection/containment trenches near the right of way.
    • Photos of recent progress at the clean-up site have been added to the Photos and Media page.

    Safety update:

    • After six weeks of mobilization, construction, containment and clean-up operations, and more than 125,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents, and continues to have a sharp focus on safety.

    Plains has added an RSS feed to the Clean-up Updates page of the Rainbow pipeline incident webpage, and we encourage interested stakeholders to subscribe to the RSS feed to remain updated on the cleanup progress. Once you’ve subscribed to the RSS feed, you’ll receive the regular updates at the same time as they are added to the website.

  • Thu, 9 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 134 workers yesterday.
    • As of 7:00 a.m. on June 9th, 4329 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2131 m3 of clean oil (approximately 47% of the original estimated 4500 m3 released).
    • Access: Workers are preparing to move one of the rig mat roads, to accommodate the soil stockpiles and allow the previous road location to undergo soil remediation.
    • Oil recovery: Workers continue to cut and rake dried seasonal vegetation and wash the pond's shoreline and dormant beaver runs with fresh water in order to facilitate oil recovery.
    • Workers continue using vacuum trucks to collect oil from the perimeter trenches, bell holes, and collection/containment trenches near the right of way.

    Environmental update:

    • Spill site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.

    The next daily update will be issued tomorrow afternoon. Regular updates on our clean-up operations can be found on Plains Midstream Canada's Rainbow Pipeline Incident webpage at www.plainsmidstream.com.

    Plains has added an RSS feed to this Clean-up Updates page. We encourage you to subscribe to the RSS feed to remain updated on the cleanup progress. Once you've subscribed to the RSS feed, you'll receive the regular updates at the same time as they are added to the website.

  • Wed, 8 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 153 workers yesterday.
    • As of 7:00 a.m. on June 8th, 4238 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2125 m3 of clean oil (approximately 47% of the original estimated 4500 m3 released).
    • Oil recovery: There's been significant progress in removing surface oil from the ponds; four skimmers are currently in operation on the pond.
    • Workers continue using vacuum trucks collect oil from the perimeter trenches, bell holes, and collection/containment trenches near the right of way.
    • Remediation continues on the top layer of contaminated soil, as it is removed and stockpiled on the north end of the right of way.

    Plains has added an RSS feed to this Clean-up Updates page. We encourage you to subscribe to the RSS feed to remain updated on the cleanup progress. Once you've subscribed to the RSS feed, you'll receive the regular updates at the same time as they are added to the website.

  • Tue, 7 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 146 workers yesterday.
    • As of 7:00 a.m. on June 7th, 4057 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2114 m3 of clean oil (approximately 47% of the original estimated 4500 m3 released).
    • Oil recovery: Workers continue to wash the pond’s shoreline and flush dormant beaver runs with fresh water in order to facilitate oil recovery. Debris removal continues in the pond areas.
    • Soil remediation: Contaminated soil continues to be removed and stockpiled on the north end of the right of way.

    Environmental update:

    • Spill site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.
  • Mon, 6 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 139 workers over the weekend.
    • As of 7:00 a.m. on June 6th, 3871 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2108 m3 of clean oil (approximately 47% of the original estimated 4,500 m3 released). Cold, snowy weather and rain has affected recovery efforts.
    • Oil recovery: Workers continue to wash the pond’s shoreline with fresh water and cut and rake dried seasonal vegetation in order to facilitate oil recovery. Vacuum trucks continue to be used to remove collected oil from containment trenches and bell holes.
    • Over the weekend, an average of 5 skimmers were in operation in the pond areas daily. There’s been significant progress in removing surface oil from the ponds, so fewer skimmers are required for the remaining area.
    • Soil remediation: Remediation continues on the top layer of contaminated soil, as it is removed and stockpiled on the north end of the right of way.
  • Fri, 3 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 166 workers yesterday.
    • As of 7:00 a.m. on June 3rd, 3563 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2079 m3 of clean oil (approximately 46% of the original estimated 4,500 m3 released).
    • Oil recovery and remediation: Work continues on clean up in the pond area. Workers are using a boat and wearing chest waders, so they are able to reposition booms that corral oil to skimmers for recovery.
    • Rig mats and collection/containment trenches have been installed down the right of way to create access for oil recovery operations. Workers use vacuum trucks to also collect oil from the collection/containment trenches excavated around the stockpiles of impacted soil.
    • Photos of recent progress at the clean-up site have been added to the Photos and Media page.

    Safety Update:

    • After 35 days of mobilization, construction, containment and clean-up operations, and more than 107,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents, and continues to have a sharp focus on safety. Daily safety briefings focus on recognizing fatigue, identifying potential hazards on site, reiterating known site hazards, and general safety awareness.
  • Thu, 2 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 137 workers yesterday.
    • As of 7:00 a.m. on June 1st, 3440 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2064 m3 of clean oil (approximately 46% of the original estimated 4,500 m3 released).
    • Oil recovery and remediation: Workers continue clean-up efforts on the pond area, using canoes and wearing hip waders to remove oily woody debris and vegetation
    • Rig mats have been placed at strategic points around the pond area to allow for quick deployment of skimmers to be optimally positioned according to weather conditions and wind direction.

    Safety Update:

    • Site safety continues to be a key focus for the clean-up operations. Daily safety briefings continue to be conducted to ensure workers understand precautions for working in the heat, are aware of how to work safely around power lines, and to maintain constant attention to safe operations.
  • Wed, 1 Jun 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 151 workers yesterday.
    • As of 7:00 a.m. on June 1st, 3271 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2047 m3 of clean oil (approximately 45% of the original estimated 4,500 m3 released).
    • Oil recovery and remediation: Workers continue flushing the pond’s shoreline with fresh water to remove as much oil as possible, and adjusting booms to corral remaining product in order to facilitate recovery.
    • Volumes continue to be recovered from bell holes and the containment trenches along the perimeter of the site.
    • Remediation of the treed area adjacent to the right of way continues, as trees are removed to allow the layer of contaminated soil to also be removed. The felled trees will be used in future reclamation efforts in the area, once the remediation process has been completed.

    Environmental Update:

    • Spill site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.  

     

  • Tue, 31 May 2011 13:30:00 MST

    Site update:

    • Daily personnel count on site peaked at 145 workers yesterday.
    • As of 7:00 a.m. on May 31st, 3197 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2035 m3 of clean oil (approximately 45% of the original estimated 4,500 m3 released).
    • Access: Perimeter mats continue to be repaired and maintained. Additional mats are being placed to allow greater access throughout the clean-up area.
    • Oil recovery and remediation: Workers are washing the pond’s shoreline with fresh water and cutting and raking dried seasonal vegetation to remove as much oil as possible.
    • Remediation continues on the top layer of contaminated soil, as it is removed and stockpiled on the north end of the right of way.
    • Tree bunchers are being used to remove trees from the north side of the right of way. Once the trees have been removed, the layer of contaminated soil will be ‘peeled back’ from the area for safe handling. The felled trees will be used in future reclamation efforts in the area, once the remediation process has been completed.

     

  • Mon, 30 May 2011 13:30:00 MST

    Site Update:

    • Daily personnel count on site peaked at 140 workers over the weekend.
    • As of 7:00 a.m. on May 30th, 3,087 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 2,007 m3 of clean oil (approximately 45% of the original estimated 4,500 m3 released).
    • Staff continue to focus their oil recovery efforts on the pond area, removing oily woody debris and vegetation, as well as using large fans to move surface oil to skimmers.
    • 13 skimmers were in operation yesterday, with 12 vacuum trucks working on-site.

     

  • Fri, 27 May 2011 13:30:00 MST

    Operations Update:

    • Since our initial spill response and containment efforts, work on the site has shifted from initial construction activities to improve access for product recovery work, to a heightened focus on oil recovery and clean up, with continued activities to minimize wildlife and environment impacts.
    • Daily personnel count on site remains at about 143 workers.
    • Oil recovery continues, using vacuum trucks and skimmers. Crews have been repositioning skimmers to maximize oil recovery.
    • As of 7:00 a.m. on May 27th, 2,671 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 1,887 m3 of clean oil (approximately 42% of the original estimated 4,500 m3 released).
    • Remediation activities to remove and stockpile the top layer of contaminated soil on the north end of the right of way have begun.

    Environmental Update:

    • Spill site air monitoring continues, with air monitoring units stationed downwind of clean-up operations. All levels remain well within acceptable industry standards.

    Safety update:

    • After 28 days of mobilization, construction, containment and clean-up operations, and more than 95,000 staff hours logged, the Rainbow pipeline clean-up site has had no lost-time incidents, and continues to have a sharp focus on safety.  Daily safety briefings focus on recognizing fatigue, identifying potential hazards on site, reiterating known site hazards, and general safety awareness.

    The next daily update will be issued Monday afternoon with a summary of progress over the weekend.

  • Thu, 26 May 2011 13:30:00 MST

    Operations Update:

    • Daily personnel count on site remains at about 145 workers.
    • Crews continue to add and maintain rig mats throughout the site and around the perimeter. Rig mats were cleaned to ensure they remain in good repair and to reduce dust in the area.
    • Oil recovery remains a focus, with 14 skimmers in operation yesterday and 14 vacuum trucks on-site.
    • As of 7:00 a.m. on May 26th, 2,580 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 1,853 m3 of clean oil (approximately 41% of the original estimated 4,500 m3 released).  
    • Additional bell holes were excavated for further free oil collection.
    • Product continues to be recovered from containment trenches along the perimeter of the spill site.

    Environmental Update:

    • Wildlife flagging was restrung to serve as a deterrent for local wildlife.
    • One bear was sighted outside of the perimeter fencing.
  • Wed, 25 May 2011 13:00:00 MST

    Operations Update:

    • Full-scale operations resumed today with approximately 140 personnel on site.
    • Crews repaired rig mats on the perimeter and placed additional timber rig mats on the east entrance road to allow for vehicle turnarounds.
    • A dust-control truck was mobilized to suppress dust on the roads; road maintenance work has also been scheduled.
    • Product continues to be removed from containment trenches along the perimeter of the release site.
    • As of 7:00 a.m. on May 25th, 2,522 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 1,832 m3 of clean oil (approximately 41% of the original estimated 4,500 m3 released).
    • There are currently no wildfire threats to the work site.  We remain in regular contact with Alberta Sustainable Resource Development to monitor the situation.

    Environmental Update:

    • Wildlife fences were checked for damage and repaired as required.

    Pipeline Update:

    • The 24" Rainbow pipeline from Nipisi to Edmonton was restarted today at approximately 1:00 p.m.  The pipeline was shut down on May 15th as a precautionary measure due to area forest fires.
    • The 20” Rainbow pipeline from Zama Lake to Nipisi remains closed, pending restart approval by the ERCB.
  • Tue, 24 May 2011 17:00:00 MST

    Site Update:

    • Approximately 50 personnel were working on-site today to prepare for a safe return to full clean-up operations.
    • As of 7:00 a.m. on May 24, 2011, 2,433 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 1,809 m3 of clean oil (40% of the original estimated 4,500 m3 released).
    • Full-scale clean-up operations are scheduled to commence on Wednesday, May 25. We’ll shift to one daily 12-hour shift, as operations transition from construction activities to focus on maximizing our clean-up operations. 24-hour security and wildlife patrols will continue.
    • We remain in daily contact with Alberta Sustainable Resource Development to understand area wildfire activity and its implications on the work taking place at the clean-up site.

    Pipeline Update:

    • Southbound crude oil flows on the 24" Rainbow pipeline from Nipisi to Edmonton remain closed. With power recently restored, we're working with area producers to ensure a smooth return to operations.
    • The 20" Rainbow pipeline from Zama to Nipisi remains closed, pending approval by the ERCB.
  • Mon, 23 May 2011 17:30:00 MST

    Pipeline Update:

    • Southbound crude oil flows on the 24” Rainbow pipeline from Nipisi to Edmonton remain closed. We expect the line to be operational as soon as tomorrow.

    Site Update:

    • The Northern Sunrise County Emergency Operations Centre (EOC) lifted its evacuation order yesterday afternoon at the Plains Midstream Canada Rainbow pipeline release clean-up site.
    • A small crew is readying the site for a safe return to full clean-up operations; current activities include removing fallen trees and debris at the site, due to high winds.  
    • On-site work will continue to scale up tomorrow, and we anticipate resuming full-scale clean-up operations at 7:00 a.m. on Wednesday, May 25, provided site conditions haven’t changed.
    • Even with the evacuation order lifted, conditions can change quickly. We will remain in daily contact with Alberta Sustainable Resource Development to understand area wildfire activity and its implications on the work taking place at the clean-up site.

     

  • Fri, 20 May 2011 17:00:00 MST

    Pipeline update:

    • Southbound crude oil flows on the 24" Rainbow pipeline from Nipisi to Edmonton remain closed due to electrical outages related to forest fires in the area. We have been advised that it could take up to two weeks (or longer) to restore service to the area.
    • The 20” Rainbow pipeline from Zama to Nipisi remains closed, pending restart approval by the ERCB.

    Operations report:

    • Clean-up operations at Plains Midstream Canada’s Rainbow pipeline release site remain suspended due to a forest fire threat. The site remains under an evacuation order.
    • We continue to monitor the forest fire situation through regular contact with the Northern Sunrise County Emergency Operations Centre (EOC) and Alberta Sustainable Resource Development.
    • We anticipate initiating clean-up operations at 7 a.m. on Tuesday, May 24, pending the evacuation order is lifted with enough time to re-mobilize the workforce.

    Plains remains committed to keeping our stakeholders informed of the progress of our clean-up operations. As such, the next daily update will be available once clean-up operations resume.

  • Thu, 19 May 2011 17:00:00 MST

    • Clean-up operations at Plains Midstream Canada’s Rainbow pipeline release site remain suspended due to a forest fire threat. The site remains under an evacuation order.
    • We continue to monitor the forest fire situation through regular contact with the Northern Sunrise County Emergency Operations Centre (EOC) and Alberta Sustainable Resource Development.
    • Once we receive approval from the EOC, we will return to the site and begin efforts to re-mobilize the workforce.
    • We received approval from the EOC to visit the site yesterday afternoon to conduct a brief visit to assess the site and examine the containment status. The site was inspected and the oil remains contained.
    • We anticipate resuming clean-up operations at 7 a.m. on Tuesday, May 24, pending appropriate conditions.
  • Wed, 18 May 2011 17:00:00 MST

    • Clean-up operations at Plains Midstream Canada’s Rainbow pipeline release site remain suspended due to a forest fire threat.  
    • We continue to monitor the forest fire situation through regular contact with the Northern Sunrise County Emergency Operations Centre and Alberta Sustainable Resource Development.
    • Once we receive approval from the Northern Sunrise County Emergency Operations Centre, we will return to assess the site and begin efforts to re-mobilize the workforce.
    • We anticipate resuming clean-up operations on Tuesday, May 24, pending appropriate conditions.
  • Tue, 17 May 2011 17:00:00 MST

    • Due to the continued evacuation order from Northern Sunrise County, Plains Midstream Canada’s clean-up operations remain suspended.
    • The 24” southbound Rainbow pipeline from the Nipisi Terminal to Edmonton remains shut down, and will resume operations once the full impact of the forest fires is known, the evacuation order is lifted and our staff can safely return to operations in the area. 
  • Mon, 16 May 2011 16:00:00 MST

    • Plains Midstream Canada’s Rainbow pipeline clean-up operations remain suspended, as Northern Sunrise County’s evacuation order remains in effect.
    • Plains continues regular communication with the Northern Sunrise Country Emergency Operations Centre and Alberta Sustainable Resource Development as they respond to the forest fire situation.
    • The 24” southbound Rainbow pipeline remains shut down, and will resume operations once the full impact of the forest fires is known, the evacuation order is lifted, and our staff can safely return to operations in the area.  
    • Clean-up operations will resume once the evacuation order is lifted, and it is determined safe to do so. At that time, we’ll assess the site, and re-mobilize the workforce. The safety of our workers is, and will continue to be, our primary concern.
  • Sun, 15 May 2011 17:00:00 MST

    Site update

    • Clean-up operations at Plains Midstream Canada’s Rainbow pipeline release site have been suspended due to a forest fire threat.  
    • The Northern Sunrise County Emergency Response Centre issued an evacuation order effective 11:00 a.m., and all staff were safely evacuated from the clean-up site.
    • Our focus has been the safe removal of our staff from the site. We’ll continue to monitor the forest fire situation through regular contact with the Northern Sunrise County Emergency Response Centre and Alberta Sustainable Resource Development.
    • As a precautionary measure, Plains has shut down the southbound crude oil flows on the 24” Rainbow pipeline from the Nipisi Terminal to Edmonton.
    • Clean-up operations will resume once the evacuation order is lifted, and it is determined safe to do so. The safety of our workers is, and will continue to be, our primary concern.

    Operations Report  

    • Due to the site evacuation order, clean-up operations at the Rainbow pipeline site have been suspended.
    • As of 7:00 a.m., 2,272 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, including 1,759 m3 of clean oil (39% of the original estimated 4,500 m3 released).
  • Sat, 14 May 2011 17:00:00 MST

    Operations report

    • 13 skimmers in operation. Skimming work is moving toward the north end of the ponds.
    • Additional lengths of boom have been added to corral remaining product.
    • Rig mats have been installed around the full perimeter of the cleanup site; work continues on additional right-of-way matting.
    • Maintenance on the access road has improved the shoulder of the road.
    • A third air-monitoring van has arrived onsite and will be deployed as needed around the work site.
    • As of 07:00 am, 2,145 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, representing 1,680 m3 of clean oil (37% of the original estimated 4,500 m3 released).

    Environmental report

    • High winds are expected tonight, which could affect operations.
    • Small portion of the west wildlife fence was damaged by wildlife last night, and repaired today.  
    • Wildlife impact to date includes: 7 beaver, 21 duck, 13 migratory birds, 2 frogs.
    • A brush fire has been reported about 10 km away. On-site operations remain in contact with Alberta Sustainable Resource Development to monitor the situation.
  • Fri, 13 May 2011 17:00:00 MST

    Safety Update

    • After 14 days of mobilization, construction, containment and clean-up operations, and more than 78,000 staff hours logged, the Rainbow pipeline cleanup site has had no lost-time incidents, and continues to have a sharp focus on safety.  Daily safety briefings focus on recognizing fatigue, identifying potential hazards on site, reiterating known site hazards, and general safety awareness.

    Environmental Update

    • Alberta Environment's Mobile Air Monitoring Lab has left Little Buffalo School as of 11 a.m. today and is heading to Edmonton. Plains Midstream Canada has adequate air monitoring resources in place long-term at the release site and at Little Buffalo and Marten Lake schools. These private air monitoring resources have been calibrated by Alberta Environment. To date, no ambient air issues have been detected and all readings are well below Alberta's Ambient Air Quality Guidelines. All air quality data is being forwarded to Alberta Health Services and Alberta Environment for review.

    Operations report

    • Night shift resumed last night - focused on site work rather than recovery efforts due to low temperatures; third shift will continue tonight.
    • 12 skimmers in operation (2 down for routine maintenance, 1 down for repair); skimmers repositioned to maximize recovery due to wind changes.
    • Began isolating portions of the pond with booms to prevent oil movement.
    • Silt fence completed on inverted weir; water is starting to move through the weir from natural runoff.
    • 120 metres of matting is being installed, which will complete the west perimeter access road around the site.
    • 13 vacuum trucks are on duty during the day and 6 vacuum trucks are on duty at night; 1 hydro-vac is on standby.
    • As of 07:00 am, 2,016 m3 of spill fluid (which includes oil, water and solids) has been recovered from the release site, representing 1,623 m3 of clean oil (36% of the original estimated 4,500 m3 released).

    Environmental report

    • Wildlife impact to date includes: 7 beaver, 19 duck, 6 migratory birds, 2 frogs.
    • Amphibian traps are being monitored around the site.
    • Volumes continue to be recovered from the bell holes along the east and west sides of the right of way, and the containment trenches around the perimeter of the spill site.
  • Thu, 12 May 2011 17:00:00 MST

    Operations report

    • Cleanup operations resumed as scheduled at 07:00 am.
    • 14 additional staff are being brought on site to resume skimming operations from 23:00 to 08:00, providing 24 hours of recovery efforts.
    • As of 07:00 am, 1811 m3 (~1,811,000 litres) of spill fluid (which includes oil, water and solids) has been recovered from the release site, representing 1489 m3 of clean oil (33% of the original estimated 4,500 m3 released). Please note that previous figures included a small amount of water content.
    • Brushing and mulching work are ongoing, and timber rig mats continue to be placed to provide a more stable workspace to clean up the site.
    • As a result of yesterday’s work suspension at the clean-up site due to an anonymous distress call, radio frequencies were reprogrammed into all on-site radios.
    • An inverted weir was installed on the pond as a precaution in the case of heavy rains. The weir will permit water overflow while ensuring oil containment.
    • Silt fencing will be installed downstream of the weir.

    Environmental report

    • Bears were observed well outside of the fenced-off spill area last night.
    • High winds yesterday caused trees to fall in some areas of the work site. Crews are working to flag leaning trees for removal.
    • Wildlife impact to date includes: 7 beaver, 19 duck, 6 migratory birds, 2 frogs
    • Volumes continue to be recovered from the bell holes along east and west sides of the right of way, and the containment trenches around the perimeter of the spill site.

     

  • Wed, 11 May 2011 18:00:00 MST

    Operations report

    • Bell holes are being excavated along the east and west sides of the right of way for additional free oil collection.
    • As the temperature warms, the skimmers are able to operate longer hours. Preparations are being made to return to 24-hour clean-up operations, as warmer night temperatures allow.  (Temperatures below 4°C cause the oil to become too thick for the skimmers to operate).

    Environmental report

    • Wildlife impact to date includes: 7 beaver, 18 duck, and 4 migratory birds.
    • 24-hour security and wildlife patrols continue: One bear was observed outside the perimeter fencing on an access road last night.

    Communications

    • The Rainbow Incident webpage has been updated with additional Q&As and new photos of clean-up activities on site.
    • Daily updates continue to be sent to First Nations and political leaders, municipal leaders, regulatory agencies, additional stakeholders, and media.

     

  • Tue, 10 May 2011 16:30:00 MST

    Operations report

    • Skimmers continue to recover product, as temperatures permit.
    • An additional 150 metres of timber rig mats was placed on the east side of the pipeline right of way, providing additional access for clean-up efforts.
    • To date, 1599 m3 ( ~1,600,000 litres) of crude oil (~36% per cent) of the original estimated 4,500m3 released has been recovered.
    • All recovered product has been sent to an oil cleaning and treatment facility.
    •  Secondary containment has been installed to contain any migrating oil. Only small amounts of oil in the perimeter trenches have been observed and collected to date.
    • Daily personnel count on site remains at about 300 workers.
    • Currently installing an inverted weir to ensure oil containment.

     Environmental report

    • Wildlife impact to date includes: 7 beaver, 18 duck, 1 shoreline bird, 2 sparrow and 1 migratory bird.

    Communications

    • Invitations to tour the site have been forwarded to media.
  • Mon, 9 May 2011 16:30:00 MST

    Operations report

    • 24-hour security and wildlife patrols continue, complemented by two, nine-hour clean-up shifts (with one hr overlap).
    • Approximately 300 response personnel have been onsite today.
    • Approximately 30% of initial estimated crude oil release has been recovered.
    • More brushing completed on west side of pipeline right of way.
    • Placement of timber rig mats continues on the east side of the right of way
    • Recovery work continues on site, with skimming showing improved results as the temperatures increase.
    • Alberta Environment, the Alberta Energy Resources Conservation Board (ERCB), and other regulatory agencies staff remain on site.

    Environmental report

    • Wildlife impact to date includes: 7 beaver, 13 ducks, and 1 migratory bird.
    • Fish & Wildlife Management division of Alberta Sustainable Resource Development continues to wildlife protection efforts.

    Communications

    • Media site tour invitations to be sent tomorrow.
    • Rainbow Incident update website to be transitioned from Plains All American’s site to www.plainsmidstream.com tomorrow, with photos being added throughout the week.
       
  • Mon, 9 May 2011 15:00:00 MST

    Shift Change

    • Commencing today, Monday, May 9, we will be reducing the clean-up hours of on-site activities from 24-hrs per day to 17-hour days (two 9-hour shifts with one hour overlap). This change is being made for the following reasons:
      1. Weather/temperature: During the late evening and early morning, skimming operations have been hampered by dropping ambient temperatures. (Temperatures below 4 degrees Celsius cause the oil to become too thick for the skimmers to operate).
      2. Progress on site: Site construction and mobilization activities are nearing completion with the exception of brushing and clearing activities that do not operate through the evening hours.
      3. Safety: Safety considerations of contract operators and Plains staff due to fatigue.
    • It is expected that the shift change will increase onsite productivity and reduce the risk of a safety-related incident. On-site security and wildlife patrols will remain to operate on a 24-hour schedule.
    • As evening temperatures rise to allow for efficient skimming operations through the evening hours, a third shift will be implemented to resume 24-hour operations.
    • PMC will develop a mobilization plan for the third shift in the next five days to ensure the clean-up operation can be returned to 24/7 as the overnight temperatures improve.
    • PMC has secured numerous skimmers and vacuum trucks with additional equipment en route. Because of this, it is not expected that the pace of oil recovery will be impacted by the shift change.
    • Plains is committed to maintaining the safety of its employees and contractors and is working to ensure activities are carried out in a safe and efficient manner.
  • Sun, 8 May 2011 18:00:00 MST

    Operations Report

    • The release remains contained to a relatively small footprint by natural topography
    • Work continues 24 hours a day with over 300 response personnel onsite. Operating 15 skimmers with one additional skimmer to be set up by the end of day today.
    • To date, 1,326 m3 (1,350,000 liters) of crude oil (29 per cent) of the original estimated 4,500 m3 released has been recovered.
    • All recovered crude oil has been taken to offsite disposal.
    • 150 m of timber rig mats placed on east side.
    • Both East and West perimeter secondary containment trench complete.
    • 150 m of brush clearing complete on west side of the pipeline right of way
    • Four loads of timber rig mats arrived last night and will be placed on the west side.

    Environmental Report

    • The wildlife perimeter fence is now complete.
    • Continue to maintain active and passive measures of wildlife deterrents including: scare crows, bright colored flags, audio bird deterrents, mylar ribbons, effigies, and foot patrols.
    • Wildlife impact to date includes: 7 beaver, 11 ducks and 1 migratory bird.
    • Air monitoring continues at the site. No change to record of meeting Alberta Ambient air quality guidelines. Air monitoring continues at the school at Little Buffalo. No change to zero level readings for hydrocarbons.

    Communications

    • Plan to provide high-resolution pictures of the site available by mid-week.
    • Invitations for site tours to be sent tomorrow.

    Pipeline

    • Plains Midstream Canada has determined that the cause of the leak was not due to corrosion or other age-dependent cause; it was most likely due to a settlement/compaction issue resulting from an excavation in 2010.
    • We are confident the pipeline is ready and safe to operate and can resume operation independently from the clean-up operation.
    • Have completed repairs to pipeline.
    • Have submitted application for restart.
    • ERCB is conducting its own investigation as to cause.
    • Line restart contingent on receiving regulatory approval.
  • Sat, 7 May 2011 22:00:00 MST

    • 9 skimmers operating (2 added this afternoon). 5 more skimmers expected tomorrow
    • Perimeter fence 100% complete
    • 25% of estimated oil release has been recovered. Total recovered: ~1166 m3; 7,345 barrels

     

  • Wed, 4 May 2011 14:05:00 MST

  • Wed, 4 May 2011 14:00:00 MST

    Plains has monitored air quality at the site of the crude oil pipeline release since day one of the release. The monitors at the spill site have detected no hydrocarbon levels above Alberta Ambient Air Quality guidelines. Additionally, immediately upon hearing of concerns raised at the school in the community of Little Buffalo, approximately 14 KM from the spill, the company began air quality monitoring at that location. These monitors have been in place since Monday, and have detected no hydrocarbon levels whatsoever. We will continue monitoring air quality at the spill site and in surrounding areas.

  • Wed, 4 May 2011 12:30:00 MST

    We have recovered approximately 300 m3 (~1900 bbls) of the oil (approximately 7% of estimated spill volume). Recovery efforts continue 24 hrs per day.

  • Wed, 4 May 2011 08:00:00 MST

    Physical pipeline repairs have been completed. We are awaiting regulatory approval to restart the line. The oil has been contained and recovery efforts are ongoing.

  • Wed, 4 May 2011 01:00:00 MST

    The pipeline has been excavated and examined by the Alberta authorities. Based on this inspection and the recently interpreted results of pipeline integrity tool runs, it appears this is a singular failure and not a systemic problem. Plains is working to repair the line and bring it into service as quickly as possible.

  • Tue, 3 May 2011 23:30:00 MST

    Plains Midstream Canada (PMC) is fully engaged in an aggressive response to a pipeline spill in northern Alberta on the company's Rainbow Pipeline. The oil is physically contained within a 0.8 kilometers (half a mile) by 50 meter (150 feet) area.

    Facts about the emergency response:

    • Since day one, PMC has engaged in a 24-hour response effort.
    • PMC Sr. Management was involved from the onset and arrived on the scene within 12 hrs to lead the response effort.
    • The spill site is located in a very remote and densely forested area, requiring a substantial construction effort in connection with oil recovery efforts.
    • Two roads were built through the forest so workers and equipment could be transported to the north and south ends of the spill site.
    • PMC secured 1400 wooden mats ("rig matting") from across Alberta to build temporary wooden access roads over the soft muskeg conditions.
    • Approximately 100 response personnel are currently on the scene and involved in recovery and remediation operations.
    • Specialty all-terrain heavy equipment were on site since day two, including 10 amphibious track-mounted vehicles for transporting workers and material.
    • Three transfer stations with tanks and trucking facilities have been constructed to facilitate the transfer of collected oil and servicing of vehicles.
    • Electrical generators are on site and 25 lighting plants are being used so work can continue through the night.
    • 46 pieces of heavy equipment have been brought to the site to support the cleanup.

    "We have a full complement of specialists at the scene," stated Roy Lamoreaux, Director Investor Relations. "We have secured all necessary people and equipment for our response and cleanup effort."

    "The oil is in place, and we expect to retrieve a substantial portion of this spill with our recovery equipment. A remediation program will follow the cleanup, which will include removal and replacement of contaminated soil and replanting vegetation on the affected area," said Lamoreaux.

    PMC is working closely with a number of regulatory agencies including Alberta Environment and the Alberta Energy Resources Conservation Board "ERCB," and providing reports to other government agencies involved in the spill.

     

  • Tue, 3 May 2011 13:30:00 MST

    Recovery:

    • 3 skimmers are now deployed on site (up from 2);
    • Another 2 skimmers are on site and will be established to increase collection rate;
    • Oil collection and removal continues.

    Repair:

    • 1st stopple fitting installed and plug set isolating the rupture from the north side further isolating the rupture from the north;
    • 2nd stopple fitting welded, plug to be set Tuesday further isolating the rupture from the south;
    • Replacement pipe has been hydrotested.
    • 3rd party lab has been appointed and protocols established for analysis of failed pipe section (to be sent to lab post inspection by ERCB);
    • Target to have physical line repair completed by end of this week;
    • Actual restart timing contingent on regulatory approval.

    Other:

    • Investigation into cause continues.

     

  • Mon, 2 May 2011 05:40:00 MST

    • Oil recovery operation continued through the night.
    • Product collected on site and shipments to Nipisi commenced at 2:50 a.m. (ahead of the 4:00 a.m. target).
  • Sun, 1 May 2011 21:30:00 MST

    • Approximately 100 response personnel on site.
    • Road access with rig matting is complete at South entrance to spill site. (The spill site is located in a wet/muddy and wooded environment, which has required substantial work to provide vehicle access. Work required to provide access includes: clearing trees and building a road of timber "rig matting" so rubber-tired vehicles can travel on the newly matted roads).
    • Oil skimmers have been established and oil collection has commenced.
    • Plains Midstream Canada's Nipisi terminal (south of affected area) configured to accept delivery of skim oil.
    • Continuing to place matting to establish road access to the repair site and north end of the spill.

     

  • Sat, 30 Apr 2011 16:20:00 MST

    • Remediation efforts are ongoing and active 24 hrs per day. The following resources have been deployed to assist in the remediation efforts;
    • 63 response personnel are now on site, another 41 are enroute or on standby nearby;
    • 46 pieces of equipment are on site including: vehicles, generator sets, office trailers, spill response trailers, medical aid equipment, etc.;
  • Sat, 30 Apr 2011 10:50:00 MST

    Approximately 50% of the daily volumes handled by Rainbow (movements from Nipisi south to Edmonton) are unaffected.